OpenStudy (anonymous):

Steve takes 6 hours to get the yard work done by himself. If Bill works alone, it takes him 10 hours. How long would it take them working together to do all of the yard work?

OpenStudy (anonymous):

10 + 6 ------- 2 8 hours

OpenStudy (amistre64):

how much work is done in an hour by each: s = 1/6 of the yard in one hour b = 1/10 of the yard in an hour working together they can do: 1/6 + 1/10 in an hour; or 16/60; we can invert this number to determine how long it takes to finish the yard 60/16 hours

OpenStudy (amistre64):

i get 3 hours and 45 minutes ...